or the most recent fiscal year, book value of long-term debt at Schlumberger was $12,240 million. The market value of this long-term debt is approximately equal to its book value. Schlumberger’s share price currently is $59.95. The company has 1,000 million shares outstanding.

Managers at Schlumberger estimate that the yield to maturity on any new bonds issued by the company will be 13.79%. Schlumberger’s marginal tax rate would be 35%.

Schlumberger’s beta is 0.72. Suppose that the expected return on the market portfolio is 8% and the risk-free rate is 2%.

Assume that the company will not change its capital structure. Also assume that the business risk of the projects under consideration is about the same as the business risk of Schlumberger as a whole.

What would Schlumberger’s after-tax WACC be, given this information?

Do not round at intermediate steps in your calculation. Express your answer in percent. Round to two decimal places. Do not type the % symbol.
Answer #1

D/E = MV of debt/(shares*price) = 12240/(59.95*1000)

=0.2042

D/A = D/(E+D)
D/A = 0.2042/(1+0.2042)
=0.1696
Weight of equity = 1-D/A
Weight of equity = 1-0.1696
W(E)=0.8304
Weight of debt = D/A
Weight of debt = 0.1696
W(D)=0.1696
Cost of equity
As per CAPM
Cost of equity = risk-free rate + beta * (expected return on the market - risk-free rate)
Cost of equity% = 2 + 0.72 * (8 - 2)
Cost of equity% = 6.32
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 13.79*(1-0.35)
= 8.9635
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=8.96*0.1696+6.32*0.8304
WACC =6.77%